OpenStudy (e.cociuba):

For each quadratic equation, find the: a.) roots b.) y-intercept c.) Vertex y=x^2-7x+12

OpenStudy (anonymous):

a) x=3 or x=4 b) if x=0, y=12 c) change this form y=ax^2 +bx+c → y= a(x-p)^2 + q vertex is (p,q) so, y=x^2-7x+12 is →y=(x-7/2)^2-1/4 and, vertex is (7/2,-1/4) maybe right..

OpenStudy (e.cociuba):

Awesome! Thanks! :))